OpenStudy (ggscupcakes443):

WILL MEDAL .at maximum speed an airplane travels 1720 miles against wind in 5 hours. Flying the wind the plane can travel the same distance in 4 hours . Let x bE the maximum speed of the plane and y be the wind speed of the plane. What is the speed of the plane with no wind

OpenStudy (mattfort3212):

430 mph

OpenStudy (ggscupcakes443):

NO

OpenStudy (_malice_):

(X-Y) 5 =1720 That is, the speed of the plane minus the speed of the wind over 5 hours gives you that distance (X+Y) 4 = 1720 Similar reasoning. (X-Y)5 = 1720 (X+Y)4 = 1720 Lets solve the first equation first. You get X - Y = 344 so X = 344 + Y So we take that and replace X on the second equation by 344 + Y (344 + Y + Y)4 = 1720 344 + 2Y = 430 2 Y = 86 Y = 43. Thats the speed of the wind. Now we plug that value of Y on the equation we got for X earlier, resulting in X = 344 + 43 = 387. So that, 387, is the velocity of the plane without the effect of the wind.

OpenStudy (_malice_):

Was that right.